# Circumstellar Interaction in the Ultraviolet Spectra of SN 2023ixf 14–66 Days After Explosion

### v1

@@ 7 Summary
We present UV observations of SN 2023ixf 14–66 days after explosion. SN 2023ixf is UV bright, with continuum emission through day 24. Its UV spectrum shows weaker metal absorption than other Type II supernovae that do not exhibit definitive signs of CSM interaction. In the optical, the $\mathrm{H\alpha}$ P Cygni profile is slow to develop and has a shallow absorption component which we infer is due to the additional emission from CSM interaction. At day 66, the Mg II profile and $\mathrm{H\alpha}$ emission component have a similar extent in velocity space, but extremely different profile shapes, resulting from the lack of UV continuum from the supernova ejecta. We examine the contribution of different species to the UV spectrum of SN 2023ixf on days 14 and 19, finding that the majority of the absorption in the UV spectrum is due to iron, nickel, and magnesium at both epochs, with additional contributions by chromium at the later epoch.
@@
We compare SN 2023ixf to CMFGEN models of supernova ejecta with power from CSM interaction (Dessart & Hillier, 2022). Impressively, even though these models were not made for a supernova like SN 2023ixf which interacted with dense CSM during its early evolution, they match the overall evolution and the presence of individual features. It is likely that the cool dense shell in SN 2023ixf is broader and centered at a lower velocity (8500 vs. 11,700 $\mathrm{km~{}s^{-1}}$) and higher temperature than the models. This is confirmed by a custom model which places the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$. Additionally, the predicted presence of the Mg II emission after 50 days is confirmed, demonstrating that SN 2023ixf is interacting with CSM at all epochs through day 66.
@@
We visually identify the model that best matches our observed spectrum at each epoch, finding that the power from CSM interaction decreases with time. From this we conclude that at larger radii, there is lower density CSM, suggesting that the mass-loss rate increased rapidly just prior to explosion. While the flash features favor a confined CSM, these disappear after a week, limiting the radius out to which the mass-loss history can be traced by this method. Our UV observations are sensitive to much lower CSM densities, and we use them to trace the CSM density to significantly larger radii and the mass-loss history to earlier epochs. With a shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$, our observations probe the CSM density between a radius of $1.2\times 10^{15}$ cm to $5.2\times 10^{15}$ cm, and show the density evolving from $5.2\times 10^{-16}$  $\mathrm{g~{}cm^{-3}}$ to $4.8\times 10^{-20}$  $\mathrm{g~{}cm^{-3}}$ over this range. Using the wind velocity of 55 $\mathrm{km~{}s^{-1}}$, we find that the mass-loss rate decreases from $8.7\times 10^{-4}$  $\mathrm{M_{\odot}~{}yr^{-1}}$ in our first observation to $1.7\times 10^{-6}$  $\mathrm{M_{\odot}~{}yr^{-1}}$ in our last observation, showing the dramatic change in mass-loss rates 33–7 yr before explosion. Additionally, the mass-loss rate derived from the day 66 spectrum shows that 33 yr before explosion, the RSG progenitor’s mass-loss rate was consistent with a quiescent RSG wind (Beasor et al., 2020).
@@
We also compare SN 2023ixf with other supernovae having UV spectra around days 10 and 20. These objects cover a range of UV fluxes, similar to the range spanned by the CMFGEN models with varying levels of CSM interaction. This suggests that we can use the UV to sensitively distinguish varying levels of CSM interaction. Around day 10, we see that Mg II is clearly present in absorption in the spectra with the most CSM interaction. Often appearing next to it is Fe II; this can be seen as a distinct feature in SN 2023ixf and SN 2022wsp, and it may be present in the other supernovae either blended with Mg II owing to its formation at higher velocities or as a weaker feature. Mg II and Fe II are weaker around day 20 in all supernovae. Additionally, the apparent emission line at 1924 Å is caused by a window of lower iron absorption, rather than emission from an individual species, demonstrating the complexity in associating a single species with an individual feature in UV spectra.
@@
Based on the shape and width of the Mg II emission in the day 66 spectrum, we conclude that this emission is from the cool dense shell which is well described by a thick shell with a Gaussian density profile: FWHM $\approx 7700$ $\mathrm{km~{}s^{-1}}$and an inner radius velocity of $\sim-6500$ $\mathrm{km~{}s^{-1}}$. We model the emission with an asymmetric Gaussian shell, finding that the profile is consistent with more of the observed emission coming from the blue side than the red side. Additionally, in other supernovae, this feature is either symmetric or asymmetric with a higher blue side, leading us to conclude that the profile shape is due to attenuation of the far side of the shell, likely from opacity in the shell and ejecta at early times and dust attenuation in the later-time spectra. (e.g. SN 1998S at 240d)
@@
Finally, we compile the density and mass-loss measurements of SN 2023ixf from the literature using a consistent shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$ and a wind velocity of 55 $\mathrm{km~{}s^{-1}}$. We show that the density profile from $9.5\times 10^{13}$ cm to $7.5\times 10^{15}$ cm (0.5–42 yr before explosion) decreases from $1.1\times 10^{-12}$ $\mathrm{g~{}cm^{-3}}$ to $4.9\times 10^{-20}$ $\mathrm{g~{}cm^{-3}}$. With a wind velocity of 55 $\mathrm{km~{}s^{-1}}$, this corresponds to a change in mass-loss rate of $\sim 10^{-2}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ to $\sim 10^{-6}$ $\mathrm{M_{\odot}~{}yr^{-1}}$. This shows that over the final $\sim 40$ yr of the progenitor’s life, the mass-loss rate went from that of a quiescent RSG to an extremely high mass-loss rate. The densities and mass-loss rates of these different techniques span about an order of magnitude, highlighting the uncertainties and systematic errors in our methods, even when some consistent assumptions are made.
@@
This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-26555. These observations are associated with programs GO-17313 and GO-17205. Huge thanks to our HST program coordinator Alison Vick and our contact scientist Dan Welty for assistance in planning and executing the observations. The data described here may be obtained from the MAST archive at https://dx.doi.org/10.17909/787k-c897 (catalog doi:10.17909/787k-c897). K.A.B. is supported by an LSST-DA Catalyst Fellowship; this publication was thus made possible through the support of grant 62192 from the John Templeton Foundation to LSST-DA. A.V.F. is grateful for financial assistance from the Christopher R. Redlich Fund and many other donors. Time domain research by D.J.S. and team is supported by NSF grants AST-1908972 and 2108032, and by the Heising-Simons Foundation under grant No. 20201864. Research by Y.D., S.V., N.M.R, E.H., and D.M. is supported by NSF grant AST-2008108.

### v2

@@ 7 Summary
We present UV observations of SN 2023ixf 14–66 days after the explosion. SN 2023ixf is UV bright, with continuum emission through day 24. Its UV spectrum shows weaker metal absorption than other Type II supernovae that do not exhibit definitive signs of CSM interaction. In the optical, the $\mathrm{H\alpha}$ P Cygni profile is slow to develop and has a shallow absorption component, which we infer is due to the additional emission from CSM interaction. On day 66, the Mg II profile and $\mathrm{H\alpha}$ emission component have a similar extent in velocity space, but extremely different profile shapes, resulting from the lack of UV continuum from the supernova ejecta. We examine the contribution of different species to the UV spectrum of SN 2023ixf on days 14 and 19, finding that the majority of the absorption in the UV spectrum is due to iron, nickel, and magnesium at both epochs, with additional contributions by chromium at the later epoch.
@@
We compare SN 2023ixf to CMFGEN models of supernova ejecta with power from CSM interaction (Dessart & Hillier, 2022). Impressively, even though these models were not made for a supernova like SN 2023ixf, which interacted with dense CSM during its early evolution, they match the overall evolution and the presence of individual features. It is likely that the cool dense shell in SN 2023ixf is broader and centered at a lower velocity (8500 vs. 11,700 $\mathrm{km~{}s^{-1}}$) and higher temperature than the models. This is confirmed by a custom model that places the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$. Additionally, the predicted presence of the Mg II emission after 50 days is confirmed, demonstrating that SN 2023ixf is interacting with CSM at all epochs through day 66.
@@
We visually identify the model that best matches our observed spectrum at each epoch, finding that the power from CSM interaction decreases with time. From this, we conclude that at larger radii, there is lower density CSM, suggesting that the mass-loss rate increased rapidly just prior to the explosion. While the flash features favor a confined CSM, these disappear after a week, limiting the radius out to which the mass-loss history can be traced by this method. Our UV observations are sensitive to much lower CSM densities, and we use them to trace the CSM density to significantly larger radii and the mass-loss history to earlier epochs. With a shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$, our observations probe the CSM density between a radius of $1.2\times 10^{15}$ cm to $5.2\times 10^{15}$ cm and show the density evolving from $5.2\times 10^{-16}$  $\mathrm{g~{}cm^{-3}}$ to $4.8\times 10^{-20}$  $\mathrm{g~{}cm^{-3}}$ over this range. Using the wind velocity of 55 $\mathrm{km~{}s^{-1}}$, we find that the mass-loss rate decreases from $8.7\times 10^{-4}$  $M\mathrm{{}_{\odot}~{}yr^{-1}}$ in our first observation to $1.7\times 10^{-6}$  $M\mathrm{{}_{\odot}~{}yr^{-1}}$ in our last observation, showing the dramatic change in mass-loss rates 33–7 yr before the explosion. Additionally, the mass-loss rate derived from the day 66 spectrum shows that 33 yr before the explosion, the RSG progenitor’s mass-loss rate was consistent with a quiescent RSG wind (Beasor et al., 2020).
@@
We also compare SN 2023ixf with other supernovae having UV spectra around days 10 and 20. These objects cover a range of UV fluxes, similar to the range spanned by the CMFGEN models with varying levels of CSM interaction. This suggests that we can use the UV to sensitively distinguish varying levels of CSM interaction. Around day 10, we see that Mg II is clearly present in absorption in the spectra with the most CSM interaction. Often appearing next to it is Fe II; this can be seen as a distinct feature in SN 2023ixf and SN 2022wsp, and it may be present in the other supernovae either blended with Mg II owing to its formation at higher velocities or as a weaker feature. Mg II and Fe II are weaker around day 20 in all supernovae. Additionally, the apparent emission line at 1924 Å is caused by a window of lower iron absorption, rather than emission from an individual species, demonstrating the complexity of associating a single species with an individual feature in UV spectra.
@@
Based on the shape and width of the Mg II emission in the day 66 spectrum, we conclude that this emission is from the cool dense shell, which is well described by a thick shell with a Gaussian density profile: FWHM $\approx 7700$ $\mathrm{km~{}s^{-1}}$and an inner radius velocity of $\sim-6500$ $\mathrm{km~{}s^{-1}}$. We model the emission with an asymmetric Gaussian shell, finding that the profile is consistent with more of the observed emission coming from the blue side than the red side. Additionally, in other supernovae, this feature is either symmetric or asymmetric with a higher blue side, leading us to conclude that the profile shape is due to attenuation of the far side of the shell, likely from opacity in the shell and ejecta at early times and dust attenuation in the later-time spectra (e.g. SN 1998S at 240d).
@@
Finally, we compile the density and mass-loss measurements of SN 2023ixf from the literature using a consistent shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$ and a wind velocity of 55 $\mathrm{km~{}s^{-1}}$. We show that the density profile from $9.5\times 10^{13}$ cm to $7.5\times 10^{15}$ cm (0.5–42 yr before explosion) decreases from $1.1\times 10^{-12}$ $\mathrm{g~{}cm^{-3}}$ to $4.9\times 10^{-20}$ $\mathrm{g~{}cm^{-3}}$. With a wind velocity of 55 $\mathrm{km~{}s^{-1}}$, this corresponds to a change in mass-loss rate of $\sim 10^{-2}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ to $\sim 10^{-6}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$. This shows that over the final $\sim 40$ yr of the progenitor’s life, the mass-loss rate went from that of a quiescent RSG to an extremely high mass-loss rate. The densities and mass-loss rates of these different techniques span about an order of magnitude, highlighting the uncertainties and systematic errors in our methods, even when some consistent assumptions are made.
@@
We thank the referee for their thoughtful comments. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-26555. These observations are associated with programs GO-17313 and GO-17205. Huge thanks to our HST program coordinator Alison Vick and our contact scientist Dan Welty for assistance in planning and executing the observations. The data described here may be obtained from the MAST archive at https://dx.doi.org/10.17909/787k-c897 (catalog doi:10.17909/787k-c897). K.A.B. is supported by an LSST-DA Catalyst Fellowship; this publication was thus made possible through the support of grant 62192 from the John Templeton Foundation to LSST-DA. A.V.F. is grateful for financial assistance from the Christopher R. Redlich Fund and many other donors. Time domain research by D.J.S. and team is supported by NSF grants AST-1908972 and 2108032, and by the Heising-Simons Foundation under grant No. 20201864. Research by Y.D., S.V., N.M.R, E.H., and D.M. is supported by NSF grant AST-2008108.